What are the concentrations of hydrogen ion and hydroxide ion in household ammonia, an aqueous solution of NH3 that has a pH of 11.50?
1. What is the concentration of hydroxide ion in household ammonia, an aqueous solution of NH3 that has a pH of 11.50? answer in M
2. What is the concentration of hydrogen ion in household ammonia, an aqueous solution of NH3 that has a pH of 11.50? answer in M
1
Expert's answer
2020-04-13T02:24:47-0400

1. pH= -lg[H+]

[H+]=10^-pH=10-11.5 = 3.16 * 10-12

2.pOH=14 - pH=14-11.5= 2.5

pOH=-lg[OH-];

[OH-]=10-2.5 = 3.16*10-3

Answer: 1. 3.16 * 10-12 M

2.3.16*10-3 M